# STHN: Deep Homography Estimation for UAV Thermal Geo-localization with Satellite Imagery

### v1

@@ III-C Two-stage Training Strategy
For training the two-stage model, we first train the coarse alignment module from scratch, and then we attach the refinement module to the end of the coarse alignment module and jointly fine-tune the two modules. We discovered that augmenting the bounding box $B$ is crucial for effectively fine-tuning the refinement module. This requirement arises because the refinement module always tends to make no or only minor adjustments if the coarse alignment already performs well on training and validation sets. Furthermore, we observed that merely fixedly expanding the cropped boxes without random shifting and enlargement does not enhance performance. To boost the refinement module’s effectiveness, we augment $B$ by shifting the center coordinates $(x_{B},y_{B})$ by $(\Delta p_{1},\Delta p_{2})$ and expanding the width $W_{B}$ by $2\Delta p_{3}$ during training. During the evaluation phase, we consistently expand $W_{B}$ by $\Delta p_{4}$ to mitigate the potential offset error of the coarse alignment.
@@ IV Experimental Setup
In this section, we outline the experimental settings, covering the dataset (Section IV-A), evaluation metrics (Section IV-B), and implementation details (Section IV-C).
@@ IV-A Dataset
For training and evaluation, our study utilizes the Boson-nighttime [11] real-world dataset which contains $10,256$ train pairs, $13,011$ validation pairs, and $26,568$ test pairs of coupled satellite RGB and nadir-view thermal imagery. We have expanded the dataset by augmenting the collection of satellite images without corresponding thermal images from $79,950$ to $163,344$ images, covering the area of $215.78~{}$\mathrm{k}\mathrm{m}$^{2}$. This enhancement focuses on the desert and farm areas near the original dataset’s sampling region, thereby incorporating a broader spectrum of geographical patterns. Additionally, the test region is then excluded from the generated data to ensure a robust evaluation of generalization performance. The thermal images in the dataset are captured between 9:00 PM and 4:00 AM, and they are aligned with an approx. spatial resolution of $1~{}$\mathrm{m}\mathrm{/}\mathrm{p}\mathrm{x}$$. The thermal images are cropped to $W_{T}\times W_{T}$ pixels ($\mathrm{p}\mathrm{x}$), where $W_{T}=512$. The satellite images  [^2] Bing RGB satellite imagery is sourced from Maxar: https://www.bing.com/maps/aerial are cropped to $W_{S}\times W_{S}$, where $W_{S}$ can be $512$, $1024$, or $1536$. Fig. 3 shows the ground truth overlap between thermal images and satellite images with different $W_{S}$. For $W_{S}=1024$ and $1536$, the overlap ratios are $25\%$ and $11\%$, if the thermal images are fully within the satellite region.
@@ IV-B Metrics
We deploy two accuracy metrics in our evaluation: Mean Average Corner Error (MACE) and Center Error (CE). MACE, extensively adopted in previous studies on deep homography estimation [13, 12, 17], measures the mean value of the average distances between the four corners of estimated and ground truth image alignments. Conversely, CE measures the mean value of the distances between the center points of predicted thermal image displacements and ground truth ones, thereby measuring geo-localization accuracy.
In our experimental analysis, the maximum spatial distance between the center points of input thermal and satellite images $D_{C}$ emerges as a critical factor influencing estimation performance. Intuitively, a larger $D_{C}$ implies a greater translation from the center required for the four-corner displacement, which in turn becomes more challenging to predict accurately. To validate the robustness of our method, we cautiously ablate results across a spectrum of $D_{C}$, demonstrating our approach’s capability to maintain reliable performance under varying degrees of challenging translations.
@@ IV-C Implementation Details
For pre-processing, the resize side length $W_{R}$ is $256~{}$\mathrm{p}\mathrm{x}$$. The training iteration numbers of the coarse alignment and refinement modules are $200000$ with a batch size of $16$. The AdamW optimizer [36] is employed for model training, utilizing a linear learning rate decay scheduler with warmup with the peak learning rate at $1\mathrm{e}{-4}$. The numbers of iterative updates $K_{1}$ and $K_{2}$ are both set to $6$. Depending on the setting, the correlation module’s level is $2$ (for $W_{S}=512$) or $4$ (for $W_{S}=1024,1536$) with a search radius of $4$. For bounding box augmentation, $\Delta p_{1},\Delta p_{2}$ is set to vary between $(-64,64)$, $\Delta p_{3}$ is set within $[0,64)$, and $\Delta p_{4}$ is $64$ by parameter tuning. Our models are developed using PyTorch. The inference speed is measured with one NVIDIA RTX-2080-Ti GPU.
@@ V Results
TABLE II: Comparison between different image-based matching methods and our homography estimation methods when $D_{C}=512$\mathrm{m}$$.
@@
In this section, we present both quantitative and qualitative analysis in Boson-nighttime [11]. For a fair comparison, we ensure the same data and pipeline consistency steps for training and evaluating across all methods. Initially, in Section V-A and V-B, we establish a baseline by assuming that thermal images are aligned to the north, facilitated by an onboard compass and a gimbal-equipped thermal camera, with our focus primarily on translation error. Subsequently, in Section V-C, we broaden our analysis to encompass geometric noises such as rotation, resizing, and perspective transformation noises.
@@ V-A Comparison with Baselines
In the results detailed in Table I, we initiate the analysis by evaluating the efficacy of traditional keypoint matching methods, such as SIFT [37], ORB [40], and BRISK [41], integrated with outlier rejection methods like RANSAC [38] and MAGSAC++ [39]. These traditional methods demonstrate a significantly high MACE alongside substantial failure rates (calculated by instances where the number of matching keypoints $\leq 10$). This underlines the challenges inherent in complex satellite-thermal alignment.
@@
Subsequently, our analysis compares our methods with various deep homography estimation frameworks, including DHN [13], LocalTrans [16], and IHN [12] (State-of-the-art method in real-time applications). These baselines are trained on the Boson-nighttime dataset with one stage for comparison. The results show the superior performance of our methods for satellite-thermal alignment and geo-localization. A notable observation from the data is the differential performance preferences across varying $D_{C}$ distances: for $D_{C}=50~{}$\mathrm{m}$$ and $D_{C}=64~{}$\mathrm{m}$$, the optimal $W_{S}$ is $512$, while for mid-range distances of $D_{C}=128~{}$\mathrm{m}$$ and $D_{C}=256~{}$\mathrm{m}$$, using $W_{S}=1024$ leads to the best results. Additionally, for the longest distance of $D_{C}=512~{}$\mathrm{m}$$, our novel two-stage method with $W_{S}=1536$ emerges as the most effective strategy. The findings indicate that for cases where $D_{C}\leq 256~{}$\mathrm{m}$$, employing our one-stage method combined with a carefully chosen $W_{S}$ emerges as the most effective strategy. We offer further explanation of the correlation between $W_{S}$ and $D_{C}$ in Section V-B.
@@
We find that our two-stage method fails to enhance performance for distances $D_{C}=50~{}$\mathrm{m}$$, $64~{}$\mathrm{m}$$, and $128~{}$\mathrm{m}$$, instead leading to a decline in accuracy. Upon examining the visualized outcomes, we observe that for smaller distances ($D_{C}\leq 128~{}$\mathrm{m}$$), the initial coarse alignment is sufficiently accurate, making the refinement module’s excessive iterative updates introduce noise into the final predictions, thereby degrading performances. Nevertheless, our two-stage approach maintains an overall MACE of less than $15~{}$\mathrm{m}$$ across all considered $D_{C}$ values, establishing robust baselines for this task. Notably, for achieving precise geo-localization at a distance of $D_{C}=512~{}$\mathrm{m}$$, this two-stage strategy demonstrates the best performance, underscoring our method’s effectiveness for large-scale search regions.
@@
We also compare with image-based solutions  (AnyLoc [29] and STGL [11]) on accuracy and latency aspects in Table II. The latency of image-based matching methods is calculated by $t_{e}\times(N_{S}+1)+t_{m}$, where $t_{e}$ is feature extraction time per image, and $N_{S}=841$ is the number of database images centered within a $1024\times 1024$ area (while the complete images cover a $1536\times 1536$ area) with a sampling stride of $35$ px following [11]. $t_{m}$ is the matching time per query. For AnyLoc, we directly apply the original DINOv2 [44] weights and fit the VLAD [45] parameters using our training data. We have observed a significant performance decline in AnyLoc, likely due to the domain gap between satellite and thermal imagery. STGL with GeM yields high accuracy but still suffers from high latency. Our methods exhibit significant enhancements in both accuracy and latency compared to these existing image-based matching techniques. Notably, our one-stage and two-stage methods achieve latency reductions to just $7.2\%$ and $13.0\%$ of the latency of STGL-GeM-ResNet50.
@@ V-C Robustness Evaluation and Visualization
In an ideal setting, the UAV onboard compass and gimbal camera would supply precise data, enabling the accurate alignment of images to the north. However, it is crucial for our algorithm to demonstrate tolerance towards certain rotation and perspective transformation inaccuracies during active flights. Additionally, understanding how our algorithm performs when there is a change in flight altitude—which results in a change of the thermal image’s coverage area, denoted as resizing noise—is essential. To assess the algorithm’s robustness under these conditions, we perform experiments that introduce specific rotation, resizing, and perspective transformation noises. For rotation disturbances, the thermal images undergo random rotations of $5^{\circ}$, $10^{\circ}$, or $30^{\circ}$. For resizing disturbances, the images are randomly scaled by a factor of $1+\Delta r$, with $\Delta r$ being either ±$\pm 0.1$, ±$\pm 0.2$, or ±$\pm 0.3$. Regarding perspective transformation disturbances, we randomly adjust the four corners of $512\times 512$ thermal images by $8~{}$\mathrm{p}\mathrm{x}$$, $16~{}$\mathrm{p}\mathrm{x}$$, or $32~{}$\mathrm{p}\mathrm{x}$$.
In Table III, we evaluate the robustness of our one-stage and two-stage strategies against a variety of geometric noise conditions with $D_{C}=512~{}$\mathrm{m}$$ and $W_{S}=1536$. The analysis indicates a significant decrease in performance for the one-stage method under these conditions, in contrast to the two-stage strategy, which demonstrates a notable robustness against geometric perturbations. Specifically, the two-stage strategy effectively maintains test MACE below $24~{}$\mathrm{m}$$ and test CE below $22~{}$\mathrm{m}$$ in most scenarios, with notable exceptions being in instances of $30^{\circ}$ rotation noise and resizing noise with a factor of $0.3$. While incremental perspective transformations have minimal impact on accuracy, large rotation and resizing noise can significantly degrade performance. This suggests the tolerance of our strategies to varying types of noise. Overall, the results validate the robustness of our methodologies against a spectrum of geometric noises, especially underscoring the two-stage strategy’s superior effectiveness and reliability in mitigating the negative effects of these disturbances. Fig. 7 further illustrates this point by showcasing visual comparisons between the failure instances of the one-stage method and the success cases of the two-stage method, thereby demonstrating the latter’s improved robustness.
@@
$5^{\circ}$ 30.58 28.90 19.12 17.44
$10^{\circ}$ 33.98 32.06 21.89 19.87
$30^{\circ}$ 51.79 49.13 35.60 33.37
@@
$8~{}$\mathrm{p}\mathrm{x}$$ 34.97 31.51 19.47 17.60
$16~{}$\mathrm{p}\mathrm{x}$$ 34.97 31.62 20.24 18.07
$32~{}$\mathrm{p}\mathrm{x}$$ 39.01 35.48 23.46 20.54
One stage
Two stages
VI Conclusions
In this paper, we presented a novel approach based on deep homography estimation for UAV thermal geo-localization tasks. We validate the capability of STHN to precisely align thermal images, captured by UAV onboard sensors, with large-scale satellite maps, achieving successful alignment even with a constant overlap of $11\%$. Additionally, we showcase STHN’s resilience to geometric distortions, which significantly enhances the reliability of geo-localization outcomes.

### v2

@@ III-C Two-stage Training Strategy
For training the two-stage model, we first train the coarse alignment module from scratch, and then we attach the refinement module to the end of the coarse alignment module and jointly fine-tune the two modules. We discovered that augmenting the bounding box $B$ is crucial for effectively fine-tuning the refinement module. This requirement arises because the refinement module always tends to make no or only minor adjustments if the coarse alignment already performs well on training and validation sets. Furthermore, we observe that merely fixedly expanding the cropped boxes without random shifting and enlargement does not enhance performance. To boost the refinement module’s effectiveness, we augment $B$ by shifting the center coordinates $(x_{B},y_{B})$ by $(\Delta p_{1},\Delta p_{2})$ and expanding the width $W_{B}$ by $2\Delta p_{3}$ during training. During the evaluation phase, we consistently expand $W_{B}$ by $\Delta p_{4}$ to mitigate the potential offset error of the coarse alignment.
@@ IV-A Dataset
For training and evaluation, our study utilizes the Boson-nighttime [11] real-world dataset which contains $10,256$ train pairs, $13,011$ validation pairs, and $26,568$ test pairs of coupled satellite RGB and nadir-view 8-bit thermal imagery. We have expanded the dataset by augmenting the collection of satellite images without corresponding thermal images from $79,950$ to $163,344$ images, covering an area of $215.78~{}$\mathrm{k}\mathrm{m}$^{2}$. This enhancement focuses on the desert and farm areas near the original dataset’s sampling region, thereby incorporating a broader spectrum of geographical patterns. Additionally, the test region is then excluded from the generated data to ensure a robust evaluation of generalization performance. The thermal images in the dataset are captured between 9:00 PM and 4:00 AM, and they are aligned with an approx. spatial resolution of $1~{}$\mathrm{m}\mathrm{/}\mathrm{p}\mathrm{x}$$. The thermal images are cropped to $W_{T}\times W_{T}$ pixels ($\mathrm{p}\mathrm{x}$), where $W_{T}=512$. The satellite images  [^1] Bing RGB satellite imagery is sourced from Maxar: https://www.bing.com/maps/aerial are cropped to $W_{S}\times W_{S}$. Fig. 3 shows the ground truth overlap between thermal images and satellite images with different $W_{S}$. For $W_{S}=512/1024/1536$, the size ratios between thermal images and satellite images are $100\%$, $25\%$, and $11\%$.
@@ IV-B Metrics
We deploy two accuracy metrics in our evaluation: Mean Average Corner Error (MACE) and Center Error (CE). MACE, extensively adopted in [13, 12, 17], measures the mean value of the average distances between the four corners of estimated and ground truth image alignments. Conversely, CE measures the mean value of the distances between the center points of predicted thermal image displacements and ground truth ones, thereby measuring geo-localization accuracy.
In our experimental analysis, the maximum spatial distance between the center points of input thermal and satellite images $D_{C}$ emerges as a critical factor influencing estimation performance. Intuitively, a larger $D_{C}$ implies a greater translation from the center required for the four-corner displacement, which in turn becomes more challenging to predict accurately. To validate the robustness of our method, we cautiously ablate results across a spectrum of $D_{C}$, demonstrating our approach’s capability under varying degrees of challenging translations.
@@ IV-C Implementation Details
For pre-processing, the resize side length $W_{R}$ is $256~{}$\mathrm{p}\mathrm{x}$$. The training iteration numbers of the coarse alignment and refinement modules are $200000$ with a batch size of $16$. The AdamW optimizer [37] is employed for model training, utilizing a linear learning rate decay scheduler with warmup with the peak learning rate at $1\mathrm{e}{-4}$. The numbers of iterative updates $K_{1}$ and $K_{2}$ are both set to $6$. Depending on the setting, the correlation module’s level is $2$ (for $W_{S}=512$) or $4$ (for $W_{S}=1024,1536$) with a search radius of $4$. For bounding box augmentation, $\Delta p_{1},\Delta p_{2}$ is set to vary between $(-64,64)$, $\Delta p_{3}$ is set within $[0,64)$, and $\Delta p_{4}$ is $64$ by parameter tuning. For geometric noises, we extend the coverage of thermal images, apply corresponding data augmentations, and center crop the thermal images to avoid black padding on their boundary. Our models are developed using PyTorch. The inference speed is measured with one NVIDIA RTX-2080-Ti GPU.
@@ V Results
Learned Keypoint Matching Methods
R2D2 [43] + RANSAC [39] 994.31 1160.73 1160.21 2400.31 2902.81 88.9%
LoFTR [43] + RANSAC [39] 1123.74 1697.33 1317.69 1269.71 2564.65 0%
@@
TABLE II: Comparison between different image-based matching methods and our estimation method when $D_{C}=512~{}$\mathrm{m}$$.
@@
In Sections V-A and V-B, we assume that thermal images are aligned to the north, facilitated by an onboard compass and a gimbaled thermal camera. Subsequently, in Section V-C, we broaden our analysis for geometric noises.
@@ V-A Comparison with Baselines
In the results detailed in Table I, we initiate the analysis by evaluating the efficacy of traditional keypoint matching methods, such as SIFT [38], ORB [41], and BRISK [42], integrated with outlier rejection methods like RANSAC [39] and MAGSAC++ [40]. We also evaluate learned keypoint methods including R2D2 [43] trained on our dataset and LoFTR [46] with pretrained weights. These methods demonstrate a significantly high MACE alongside substantial failure rates (calculated by instances where the number of matching keypoints $\leq 10$). This underlines the challenges of keypoint matching inherent in complex satellite-thermal alignment.
@@
Subsequently, our analysis compares our methods with various deep homography estimation frameworks, including DHN [13], LocalTrans [16], and IHN [12] (state-of-the-art method in real-time applications). These baselines with one-stage models are trained on the Boson-nighttime dataset. We report the baseline results considering $W_{S}=512$ as representative results since other $W_{S}$ show similar trends in our analysis. The results show the superior performance of our approach for satellite-thermal alignment and geo-localization. A notable observation from the data is the different performance preferences across varying $D_{C}$ distances: for $D_{C}=50~{}$\mathrm{m}$$ and $D_{C}=64~{}$\mathrm{m}$$, the optimal $W_{S}$ is $512$, while for mid-range distances of $D_{C}=128~{}$\mathrm{m}$$ and $D_{C}=256~{}$\mathrm{m}$$, using $W_{S}=1024$ leads to the best results. Additionally, for the longest distance of $D_{C}=512~{}$\mathrm{m}$$, our novel two-stage method with $W_{S}=1536$ emerges as the most effective strategy. The findings indicate that for cases where $D_{C}\leq 256~{}$\mathrm{m}$$, employing our one-stage method combined with a carefully chosen $W_{S}$ emerges as the most effective strategy. Further explanation of the correlation between $W_{S}$ and $D_{C}$ is in Section V-B.
@@
We find that our two-stage method fails to enhance performance for distances $D_{C}=50~{}$\mathrm{m}$$, $64~{}$\mathrm{m}$$, and $128~{}$\mathrm{m}$$, instead leading to a decline in accuracy. Upon examining the visualized outcomes, we observe that for smaller distances ($D_{C}\leq 128~{}$\mathrm{m}$$), the initial coarse alignment is sufficiently accurate, making the refinement module’s excessive iterative updates introduce noise into the final predictions, thereby degrading performances. Nevertheless, our two-stage approach maintains an overall MACE of less than $15~{}$\mathrm{m}$$ across all considered $D_{C}$, establishing robust baselines for this task. Notably, for achieving precise geo-localization at $D_{C}=512~{}$\mathrm{m}$$, this two-stage strategy demonstrates the best performance, underscoring its effectiveness for large-scale search regions.
@@
We also compare with image-based solutions  (AnyLoc [29] and STGL [11]) on accuracy and latency aspects in Table II. The latency of image-based matching methods is calculated by $t_{e}\times(N_{S}+1)+t_{m}$, where $t_{e}$ is feature extraction time per image, and $N_{S}=841$ is the number of database images centered within a $1024\times 1024$ area (while the complete images cover a $1536\times 1536$ area) with a sampling stride of $35$ px following [11], and $t_{m}$ is the matching time per query. For AnyLoc, we directly apply the original DINOv2 [47] weights and fit the VLAD [48] parameters using our training data. We observe a significant performance decline in AnyLoc, likely due to the domain gap between satellite and thermal imagery. STGL with GeM yields high accuracy but still suffers from high latency. Our method exhibits significant enhancements in both accuracy and latency compared to these existing image-based matching techniques. Notably, our one-stage and two-stage methods achieve latency reductions to just $7.2\%$ and $13.0\%$ of the latency of STGL-GeM-ResNet50.
@@ V-C Robustness Evaluation and Visualization
Ideally, the UAV onboard compass and gimbal camera would supply precise data, enabling the accurate alignment of images to the north. However, it is crucial for our algorithm to demonstrate tolerance towards certain rotation and perspective transformation inaccuracies during active flights. Additionally, understanding how our algorithm performs when there is a change in flight altitude—which results in a change of the thermal image’s coverage area, denoted as resizing noise—is essential. To assess the algorithm’s robustness under these conditions, we perform experiments that introduce specific rotation, resizing, and perspective transformation noises. For rotation disturbances, the thermal images undergo random rotations up to $5^{\circ}$, $10^{\circ}$, or $30^{\circ}$. For resizing disturbances, the images are randomly scaled by a factor of $1+\Delta r$, with $\Delta r$ varying within either ±$\pm 0.1$, ±$\pm 0.2$, or ±$\pm 0.3$. For perspective transformation, we randomly adjust the four corners of $512\times 512$ thermal images up to $8~{}$\mathrm{p}\mathrm{x}$$, $16~{}$\mathrm{p}\mathrm{x}$$, or $32~{}$\mathrm{p}\mathrm{x}$$.
In Table III, we evaluate the robustness of our one-stage and two-stage strategies against a variety of geometric noise conditions with $D_{C}=512~{}$\mathrm{m}$$ and $W_{S}=1536$. The analysis indicates a significant decrease in performance for the one-stage method under these conditions, in contrast to the two-stage strategy, which demonstrates a notable robustness against geometric perturbations. Specifically, the two-stage strategy effectively maintains test MACE below $22~{}$\mathrm{m}$$ and test CE below $20~{}$\mathrm{m}$$ in most scenarios, with notable exceptions being in instances of $30^{\circ}$ rotation noise. While incremental perspective transformations and resizing have minimal impact on accuracy, large rotation noise can significantly degrade performance. This suggests the tolerance of our strategies to different types of geometric noise. Overall, the results validate our method’s robustness and its ability to estimate these disturbances, underscoring the two-stage strategy’s superior effectiveness and reliability in mitigating the negative effects of these disturbances. Fig. 7 further illustrates this point by showcasing visual comparisons between the failure instances of the one-stage method and the success cases of the two-stage method, demonstrating the latter’s improved robustness.
@@
One stage
Two stages
VI Conclusions
This paper presents a novel deep homography estimation approach for UAV thermal geo-localization tasks. We validate the capability of STHN to precisely align thermal images, captured by UAV onboard sensors, with large-scale satellite maps, achieving successful alignment even with a size ratio of $11\%$. Additionally, we showcase STHN’s superior performances in terms of speed and accuracy with respect to several state-of-the-art approaches as well as its resilience to geometric distortions, which significantly enhances the reliability of geo-localization outcomes.
@@ VI Conclusions
$5^{\circ}$ 31.82 27.00 14.48 11.42
$10^{\circ}$ 33.78 28.40 14.55 11.90
$30^{\circ}$ 64.63 56.98 33.48 28.93
@@
$8~{}$\mathrm{p}\mathrm{x}$$ 30.53 26.41 13.70 10.94
$16~{}$\mathrm{p}\mathrm{x}$$ 34.07 27.22 16.24 11.73
$32~{}$\mathrm{p}\mathrm{x}$$ 44.03 30.35 20.90 14.11